company a pays dividend taka 8 percent share each year will not pay any dividend in next 2 years .After that it will pay taka 10 as dividend per share .That will grow at 6% in year next .and finally that will grown at a constant rate of 10 percent there after .if market interest rate 15%.

What is p0 ,p2,p4,p7


1
Expert's answer
2021-01-29T00:02:59-0500

find the formulа:

"P=\frac{D}{r}=\frac{10}{0.08}=125"

P0=P2=125

in the first two years, the interest and dividends do not change

find the formulа:

"P=\frac{D(1+g)}{i-g}=\frac{10(1+0.06)}{0.08-0.06}=\frac{10.6}{0.02}=530"

P4=530

find the formulа:

"P=\frac{D(1+g)}{i-g}=\frac{10(1+0.1)}{0.15-0.1}=\frac{11}{0.05}=220"

P7=220